OpenStudy (anonymous):

Find the smallest integer n such that 504n is a perfect square. I know it's fourteen, but could anyone please tell me a solution besides trial and error?

OpenStudy (accessdenied):

I would find the prime factorization of 504 and take the odd men out. 504 = 7*3*3*2*2*2 3*3 and 2*2 are perfect squares, 2*7 remains.

OpenStudy (anonymous):

okay so first we need to find prime factors of 504. 2^3 * 3^2 * 7=504 separate the square terms 2^2 * 3*2 *2*7 the left out are a 2 and 7 so answer is 2*7=14